# Strongly Coupled Sectors in Inflation: Gapless Theories and Unparticles

### v3

@@ 3.2 Differential Equations
In four dimensions, the isometries of de Sitter space are a subset of the four-dimensional conformal group $SO(4,2)$. Since we are considering conformal fields, we can extract extra bootstrap equations from the extra symmetries. From [Arkani-Hamed:2018kmz], the de Sitter isometries requires correlators of conformally coupled scalars exchanging a scalar operator to satisfy
@@
We follow the same logic outlined in [Arkani-Hamed:2023kig], using the techniques of twisted cohomology—in practice, integration-by-parts (IBP) and partial fractioning of the integrand—to derive the differential equation systems obeyed by $\psi$ and $\tilde{\psi}$. First, we define the singular loci of the integrands as
@@
Using the IBP technique described in [Fevola:2024nzj, Tom:2025poss], the minimal basis of $\text{d}\log$-forms for the DE system of $\psi$ is
@@
We can see that (3.56) reduces to equation (2.33) in [Arkani-Hamed:2018kmz] when $\Delta=1$, in which the exchange is a conformally coupled scalar. The differential operator $\text{O}_{2}$ can be defined easily by shifting $X_{1}\leftrightarrow X_{2}$, and we have
@@
We can bootstrap certain solutions of (3.57) completely by requiring them to have physical singularities. A detailed discussion is presented in [Westerdijk:2025ywh].
@@ 3.3 Exchange of Spinning Unparticles
In this section, we will use the scheme developed in [Arkani-Hamed:2018kmz] to derive the de Sitter four-point functions of conformally coupled scalars with spinning unparticle exchange, emphasizing the results about $S=1,\ \Delta=3$ and $S=2,\ \Delta=4$ cases, for conserved currents and the stress tensor of the unparticle.
@@
which means, based on the discussion in [Arkani-Hamed:2018kmz], it is valid to apply the same procedure to obtain the correlation functions with higher-spin unparticle exchange.
@@
and we already worked with the dimensionless combinations $u\equiv s/x$ and $v\equiv s/y$. It is shown in [Arkani-Hamed:2018kmz] that the four-point function with spin-$S$ can be written as
@@ 4.1 Weight-Shifting Operators
Until now, all the correlators that we have obtained are correlators with the external legs to be conformally coupled scalars. However the inflaton is nearly massless, so we need to weight-shift the external legs of conformally coupled scalars, as in [Arkani-Hamed:2018kmz, Baumann:2019oyu, Baumann:2020dch, Baumann:2022jpr].
@@
Since the scalar perturbation in (2.4) couples to the unparticles in a shift-symmetric way, we can directly relate the four-point functions of inflatons to conformally coupled scalars [Arkani-Hamed:2018kmz], because
@@
From [Arkani-Hamed:2018kmz, Baumann:2019oyu] we know that this implies that the four-point functions $\mathcal{F}(u,v)$ for $\phi$ and $F(u,v)$ for $\varphi$ are related in terms of
@@
This procedure offers us a chance to develop a clear insight into the analytic structure of inflationary correlators. However, since $\nabla_{\mu}\phi_{k_{1}}\nabla^{\mu}\phi_{k_{2}}$ corresponds to $(\partial_{\mu}\pi)^{2}$ in the EFT of inflation, the weight-shifting operation also restricts us to certain couplings between the Goldstone $\pi$ and the unparticles $\mathcal{O}_{\Delta}$. A more specific interacting vertex matching requires the de Sitter isometries being more strongly broken, by giving a subluminal speed of propagation to the scalar fluctuations, as discussed in [Pimentel:2022fsc]. This is beyond the scope of this work, and we will discuss more in the future directions.
@@ 4.2 Inflationary Trispectra
With respect to the spirit of [Maldacena:2002vr, Creminelli:2003iq], the four-point functions as we studied in pure de Sitter at leading order, will not feel this mild symmetry breaking, which makes it possible to calculate the late-time correlator of the curvature perturbation $\zeta$ using the correlator in de Sitter for $\phi$.
@@
To check the consistency of our result with [Green:2013rd], we consider the behavior in the collapsed limit $s\rightarrow 0$: in the collapsed limit, the leading contribution from all convergent Appell $F_{1}$ functions is $1$, which means the scaling behavior is controlled by the term in front of the Appell functions. Notice that by properly performing the integration, $\Delta\to 3/2$ is not divergent, and we can see the scaling behaviors are different in two regions:
@@
Both cases are consistent with the analysis in [Green:2013rd].
@@ 4.3 Inflationary Bispectra
However, taking the soft limit $k_{4}\rightarrow 0$ of the four-point function directly will result in zero because in the vertex the inflaton is derivatively coupled. Only if we consider a small inflaton mass which is proportional to the slow-roll parameter $\epsilon$, a non-zero result can be obtained. The associated mode function will be [Arkani-Hamed:2018kmz]
@@
The behavior of bispectra in the squeezed limit encodes important features of the underlying dynamics. In the squeezed limit, if we take $k_{3}\rightarrow 0$, then $k_{1}$ will be approximately equal to $k_{2}$. The leading behavior of $\mathcal{B}_{\phi}$ separates into two situations:
@@
    when $\Delta<3/2$, $\mathcal{B}_{\phi}$ scales for small $k_{3}$ as $k_{1}^{-3}k_{3}^{-3}\left(k_{3}/k_{1}\right)^{2\Delta-1}$;
@@
    when $\Delta\geq 3/2$, $\mathcal{B}_{\phi}$ scales for small $k_{3}$ as $k_{1}^{-3}k_{3}^{-3}\left(k_{3}/k_{1}\right)^{2}$.
The soft behavior for all $\Delta\geq\frac{3}{2}$ degenerates with the equilateral non-Gaussianity. Thus the squeezed limit is not a good detector for unparticles. In fact, for some range of complementary series particles, were we to detect them, we’d need more information to break their degeneracy with an unparticle.
@@ 5 Comments on Phenomenology
Figure 5: Left panel: Example of massive scalar exchange, $u^{-1}\tilde{F}(u,0.5)$, for four-point function of conformally coupled scalars $\varphi$ and an internal particle with $\mu=3$ [Arkani-Hamed:2018kmz]. Right panel: Example of scalar unparticle exchange $u^{-1}F(u,0.5)$, for four-point function of $\varphi$ and an internal unparticle with $\Delta=7/2$. Note that we rescaled the four-point functions by $u^{-1}$ in order to visually enhance the squeezed-limit behaviors. To illustrate the shape in the right panel, we also rescale the prefactor in (3.22) which is supposed to significantly suppress the amplitude with the increase of $\Delta$.
Figure 5 reveals a striking dichotomy between massive scalar and unparticle exchanges in the collapsed limit. The left panel exhibits characteristic oscillatory modulation for conventional massive scalar exchange. In contrast, the four-point function with unparticle exchange (right panel) manifests spectral continuity through the absence of such oscillations. This disappearance of oscillatory patterns fundamentally stems from the gapless nature of unparticles. The observed phenomenology directly parallels collider physics: where cosmological oscillations correspond to Breit-Wigner resonances in scattering experiments. Notably, introducing mass gaps together with strong coupling (at scales $O(\text{M}_{\text{Gap}})$) would restore resonant features [Strassler:2008bv].
@@ 6 Conclusions and Future Directions
From the plots presented in Section 5, three typical shapes can be identified. In general, when $\Delta<2$ and $n+1/2<\Delta<n+1$, where $n$ is a positive integer, the shapes are similar to the equilateral shape (e.g. in [Alishahiha:2004eh, Chen:2006nt, Langlois:2008wt, Arkani-Hamed:2003juy, Senatore:2004rj]). When $n<\Delta<n+1/2$, the shapes are similar to the orthogonal shape [Senatore:2009gt]. The new shapes given by $\Delta\sim n+1/2$ are somewhat “in between” the equilateral shapes and the orthogonal shapes, in which the equilateral limit is not a maximum but rather a saddle point. It would be interesting to see which values of $\Delta$ have small “cosine” compared to the standard shapes of non-Gaussianity [Babich:2004gb]. Notice that all these shapes suffer from heavy suppression that increases with power of $\Delta$.
@@
  • During our discussions, conformal symmetries play an important role for properties of unparticles. In order to enhance non-Gaussianities, it would be important to incorporate the effects of a small speed of sound, as discussed in [Pimentel:2022fsc]. How would the resulting shapes change?
  • Some recent developments in combinatorial/geometric structures contained in cosmological correlators appeared in [Arkani-Hamed:2017fdk, Arkani-Hamed:2023bsv, Arkani-Hamed:2023kig, Baumann:2024mvm, Hang:2024xas, Fevola:2024nzj, Arkani-Hamed:2024jbp]. In this work we have seen that the non-integer part of the unparticle scaling dimension can also act as the “twist parameter.” Can we systematically develop a theory of differential equations for tree-level unparticle diagrams?
  • We discussed the unparticles as a hypothetical field with certain symmetries. Can we develop some practical models for these unparticles in the strongly coupled sector? For example, the four-dimensional Banks-Zaks model [Banks:1981nn] becomes a gapless unparticle model in the IR [Georgi:2008pq], and is asymptotically free in the UV. Ideally, the models should contain parameters which could characterize the gapped/gapless phases. A recent model building on this scenario is discussed in [Yang:2025apy].
@@ Acknowledgements
Here we present some properties of the gamma function $\Gamma(z)$, hypergeometric function $_{2}F_{1}$ and Appell function $F_{1}$ that we used in our calculations. A more comprehensive introduction to generalized hypergeometric functions can be found in [Slater:1966fan].
@@
From [Slater:1966fan], we know that Appell $F_{1}$ function relates to hypergeometric function $_{2}F_{1}$ through a Mellin-Barnes integration:
@@ Appendix C Shape of Bispectra with Spinning Exchange
The induced bispectrum for spin-2 stress tensor exchange will be [Arkani-Hamed:2018kmz]:

### v4

@@ 3.2 Differential Equations
In four dimensions, the isometries of de Sitter space are a subset of the four-dimensional conformal group $SO(4,2)$. Since we are considering conformal fields, we can extract extra bootstrap equations from the extra symmetries. From [12], the de Sitter isometries requires correlators of conformally coupled scalars exchanging a scalar operator to satisfy
@@
We follow the same logic outlined in [92], using the techniques of twisted cohomology—in practice, integration-by-parts (IBP) and partial fractioning of the integrand—to derive the differential equation systems obeyed by $\psi$ and $\tilde{\psi}$. First, we define the singular loci of the integrands as
@@
Using the IBP technique described in [93, 94], the minimal basis of $\text{d}\log$-forms for the DE system of $\psi$ is
@@
We can see that (3.56) reduces to equation (2.33) in [12] when $\Delta=1$, in which the exchange is a conformally coupled scalar. The differential operator $\text{O}_{2}$ can be defined easily by shifting $X_{1}\leftrightarrow X_{2}$, and we have
@@
We can bootstrap certain solutions of (3.57) completely by requiring them to have physical singularities. A detailed discussion is presented in [90].
@@ 3.3 Exchange of Spinning Unparticles
In this section, we will use the scheme developed in [12] to derive the de Sitter four-point functions of conformally coupled scalars with spinning unparticle exchange, emphasizing the results about $S=1,\ \Delta=3$ and $S=2,\ \Delta=4$ cases, for conserved currents and the stress tensor of the unparticle.
@@
which means, based on the discussion in [12], it is valid to apply the same procedure to obtain the correlation functions with higher-spin unparticle exchange.
@@
and we already worked with the dimensionless combinations $u\equiv s/x$ and $v\equiv s/y$. It is shown in [12] that the four-point function with spin-$S$ can be written as
@@ 4.1 Weight-Shifting Operators
Until now, all the correlators that we have obtained are correlators with the external legs to be conformally coupled scalars. However the inflaton is nearly massless, so we need to weight-shift the external legs of conformally coupled scalars, as in [12, 13, 14, 95].
@@
Since the scalar perturbation in (2.4) couples to the unparticles in a shift-symmetric way, we can directly relate the four-point functions of inflatons to conformally coupled scalars [12], because
@@
From [12, 13] we know that this implies that the four-point functions $\mathcal{F}(u,v)$ for $\phi$ and $F(u,v)$ for $\varphi$ are related in terms of
@@
This procedure offers us a chance to develop a clear insight into the analytic structure of inflationary correlators. However, since $\nabla_{\mu}\phi_{k_{1}}\nabla^{\mu}\phi_{k_{2}}$ corresponds to $(\partial_{\mu}\pi)^{2}$ in the EFT of inflation, the weight-shifting operation also restricts us to certain couplings between the Goldstone $\pi$ and the unparticles $\mathcal{O}_{\Delta}$. A more specific interacting vertex matching requires the de Sitter isometries being more strongly broken, by giving a subluminal speed of propagation to the scalar fluctuations, as discussed in [96]. This is beyond the scope of this work, and we will discuss more in the future directions.
@@ 4.2 Inflationary Trispectra
With respect to the spirit of [11, 97], the four-point functions as we studied in pure de Sitter at leading order, will not feel this mild symmetry breaking, which makes it possible to calculate the late-time correlator of the curvature perturbation $\zeta$ using the correlator in de Sitter for $\phi$.
@@
To check the consistency of our result with [27], we consider the behavior in the collapsed limit $s\rightarrow 0$: in the collapsed limit, the leading contribution from all convergent Appell $F_{1}$ functions is $1$, which means the scaling behavior is controlled by the term in front of the Appell functions. Notice that by properly performing the integration, $\Delta\to 3/2$ is not divergent, and we can see the scaling behaviors are different in two regions:
@@
Both cases are consistent with the analysis in [27].
@@ 4.3 Inflationary Bispectra
However, taking the soft limit $k_{4}\rightarrow 0$ of the four-point function directly will result in zero because in the vertex the inflaton is derivatively coupled. Only if we consider a small inflaton mass which is proportional to the slow-roll parameter $\epsilon$, a non-zero result can be obtained. The associated mode function will be [12]
@@
The behavior of bispectra in the squeezed limit encodes important features of the underlying dynamics. In the squeezed limit, if we take $k_{3}\rightarrow 0$, then $k_{1}$ will be approximately equal to $k_{2}$. The leading behavior of $S(k_{1},k_{2},k_{3})$ separates into two situations:
@@
    when $\Delta<2$, $S(k_{1},k_{2},k_{3})$ scales for small $k_{3}$ as $\left(k_{3}/k_{1}\right)^{\Delta-1}$;
@@
    when $\Delta\geq 2$, $S(k_{1},k_{2},k_{3})$ scales for small $k_{3}$ as $\left(k_{3}/k_{1}\right)^{1}$.
The soft behavior for all $\Delta\geq 2$ degenerates with the equilateral non-Gaussianity. Thus the squeezed limit is not a good detector for unparticles. In fact, for some range of complementary series particles, were we to detect them, we’d need more information to break their degeneracy with an unparticle.
@@ 5 Comments on Phenomenology
Figure 5: Left panel: Example of massive scalar exchange, $u^{-1}\tilde{F}(u,0.5)$, for four-point function of conformally coupled scalars $\varphi$ and an internal particle with $\mu=3$ [12]. Right panel: Example of scalar unparticle exchange $u^{-1}F(u,0.5)$, for four-point function of $\varphi$ and an internal unparticle with $\Delta=7/2$. Note that we rescaled the four-point functions by $u^{-1}$ in order to visually enhance the squeezed-limit behaviors. To illustrate the shape in the right panel, we also rescale the prefactor in (3.22) which is supposed to significantly suppress the amplitude with the increase of $\Delta$.
Figure 5 reveals a striking dichotomy between massive scalar and unparticle exchanges in the collapsed limit. The left panel exhibits characteristic oscillatory modulation for conventional massive scalar exchange. In contrast, the four-point function with unparticle exchange (right panel) manifests spectral continuity through the absence of such oscillations. This disappearance of oscillatory patterns fundamentally stems from the gapless nature of unparticles. The observed phenomenology directly parallels collider physics: where cosmological oscillations correspond to Breit-Wigner resonances in scattering experiments. Notably, introducing mass gaps together with strong coupling (at scales $O(\text{M}_{\text{Gap}})$) would restore resonant features [98].
@@ 6 Conclusions and Future Directions
From the plots presented in Section 5, three typical shapes can be identified. In general, when $\Delta<2$ and $n+1/2<\Delta<n+1$, where $n$ is a positive integer, the shapes are similar to the equilateral shape (e.g. in [99, 100, 101, 102, 103]). When $n<\Delta<n+1/2$, the shapes are similar to the orthogonal shape [104]. The new shapes given by $\Delta\sim n+1/2$ are somewhat “in between” the equilateral shapes and the orthogonal shapes, in which the equilateral limit is not a maximum but rather a saddle point. It would be interesting to see which values of $\Delta$ have small “cosine” compared to the standard shapes of non-Gaussianity [105]. Notice that all these shapes suffer from heavy suppression that increases with power of $\Delta$.
@@
  • During our discussions, conformal symmetries play an important role for properties of unparticles. In order to enhance non-Gaussianities, it would be important to incorporate the effects of a small speed of sound, as discussed in [96]. How would the resulting shapes change?
  • Some recent developments in combinatorial/geometric structures contained in cosmological correlators appeared in [106, 107, 92, 108, 109, 93, 110]. In this work we have seen that the non-integer part of the unparticle scaling dimension can also act as the “twist parameter.” Can we systematically develop a theory of differential equations for tree-level unparticle diagrams?
  • We discussed the unparticles as a hypothetical field with certain symmetries. Can we develop some practical models for these unparticles in the strongly coupled sector? For example, the four-dimensional Banks-Zaks model [61] becomes a gapless unparticle model in the IR [64], and is asymptotically free in the UV. Ideally, the models should contain parameters which could characterize the gapped/gapless phases. A recent model building on this scenario is discussed in [111].
@@ Acknowledgements
Here we present some properties of the gamma function $\Gamma(z)$, hypergeometric function $_{2}F_{1}$ and Appell function $F_{1}$ that we used in our calculations. A more comprehensive introduction to generalized hypergeometric functions can be found in [112].
@@
From [112], we know that Appell $F_{1}$ function relates to hypergeometric function $_{2}F_{1}$ through a Mellin-Barnes integration:
@@ Appendix C Shape of Bispectra with Spinning Exchange
The induced bispectrum for spin-2 stress tensor exchange will be [12]: